OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

OpenStudy (uri):

First arrange it to: y=mx+c

OpenStudy (uri):

9y=11x+12

OpenStudy (mathstudent55):

The slope-intercept form of the equation of a line is \(y = mx + b\) where m = slope, and b = the y-intercept. All you need to do is solve your equation for y and compare it with the form to find what m and b are.

OpenStudy (anonymous):

i dont know how to do that, can you explain it to me?

OpenStudy (mathstudent55):

You need to solve the given equation for y. That means you want y by itself on the left side.

OpenStudy (mathstudent55):

Your equation is -11x + 9y = -12 First, we need to get rid of -11x from the left side, so we add 11x to both sides. -11x + 9y = -12 +11x +11x -------------------------- 9y = 11x - 12 Now that we have 9y on the left side, and we want y by itself, we divide both sides by 9: \( \dfrac{9y}{9} = \dfrac{11x}{9} - \dfrac{12}{9} \) Now we reduce the fractions on both sides to get: \( y = \dfrac{11}{9}x \ -\dfrac{4}{3} \) Now we see that \(m = \dfrac{11}{9}\). That is the slope. Also, \(b = - \dfrac{4}{3} \). That is the y-intercept.

OpenStudy (anonymous):

ohhh, that makes so much more sense. thanks!

OpenStudy (mathstudent55):

You're welcome.